calculate the theoretical oxygen to fuel ratio for acetylene?? Answer is 3.1
Answer #1

Acetylene : C2H2

Balanced Oxidation reaction of acetylene : C2H2 + 2.5 O2 ------> 2CO2 + H2O

Atomic weights of the atoms that make up Acetylene and oxygen:

  • Carbon (C) = 12.01
  • Oxygen (O) = 16
  • Hydrogen (H) = 1.008

*So 1 molecule of Acetylene has a molecular weight of: 2* 12.01 + 2 * 1.008 = 26.036 amu

*One oxygen molecule weighs: 2 * 16 = 32 amu

*The oxygen-fuel mass ratio is then: 2.5 * 32 / 1 * 26.036 = 80 / 26.036= 3.07

*So we need 3.07 kg of oxygen for every 1 kg of fuel.

*Rounded to single decimal place : we need 3.1 kg of oxygen for every 1 kg of acetylene.

#Hope this will help you!!

#Thank you!!!